OpenStudy (anonymous):

find the value of (16^3/4 x 4^5/4)/8

OpenStudy (anonymous):

16^3/4=8 16=(2^4) 4^5/4=4sqrt2 (8*4sqrt2)/8 4sqrt2 or 4(4th root of 4)

OpenStudy (anonymous):

2^3/2

OpenStudy (anonymous):

nope 4sqrt2 4^5/4= 4(4^1/4), 4^1/4=sqrt2

OpenStudy (anonymous):

it would be 2^5/2 if you look at it in terms of 2

OpenStudy (anonymous):

may be its write

OpenStudy (anonymous):

but i will check it

OpenStudy (anonymous):

\[16^{\frac{3}{4}}=\sqrt[4]{16^3}=2^3=8\] \[\frac{8}{8}=1\] leaving \[4^{\frac{5}{4}}=4\times\sqrt[4]{4}=4\sqrt{2}\]